### 29 Dec 2023 v2

# Large Language Model for Causal Decision Making

Abstract
Large Language Models (LLMs) have shown their success in language understanding and reasoning on general topics. However, their capability to inference based on user-specified structured data and knowledge in corpus-rare concepts like causal decision-making is still limited. In this work, we explore the possibility of fine-tuning an open-sourced LLM into LLM4Causal, which can identify the causal task, execute a corresponding function, and interpret its numerical results based on users’ queries and the provided dataset. Meanwhile, we propose a data generation process for more controllable GPT prompting and present two instruction-tuning datasets: (1) Causal-Retrieval-Bench for causal problem identification and input parameter extraction for causal function calling and (2) Causal-Interpret-Bench for in-context causal interpretation. With three case studies, we showed that LLM4Causal can deliver end-to-end solutions for causal problems and provide easy-to-understand answers. Numerical studies also reveal that it has a remarkable ability to identify the correct causal task given a query. 00footnotetext: *Corresponding author: Rui Song, songray@gmail.com. This work is not related to the author’s position in Amazon. 00footnotetext: +Equal Contribution
1 Introduction
Recently, Transformer-based LLMs containing billions of parameters are gaining popularity and are widely applied in fields such as education, legal services, and medicine (Kasneci et al., 2023; Chen et al., 2021). LLMs such as GPT-3 (Brown et al., 2020a), GPT-4, and LLaMA (Touvron et al., 2023) have shown impressive performance in multiple natural language processing tasks such as question-answering, common-sense reasoning, and translation (Zhao et al., 2023). These superior performances of LLMs have also motivated explorations on LLMs’ application on the causal decision-making process.
To explore LLMs’ potential for causal decision-making processes, recent studies are majorly based on LLM’s knowledge saved internally in its parameters. For example, Kasneci et al. (2023) utilizes LLM to recover pairwise causal relationships based on the metadata of the variables, instead of the data values. Long et al. (2023) further extends the approach to obtain causal graph estimates by querying LLM about the variable pairs. However, this line of approach has some limitations: First, this approach is limited to causal structure learning and cannot be applied to other causal tasks such as causal effect estimation and causal policy learning. Second, this type of approach is heavily based on internal knowledge in the LLM training corpus, which can be incomplete and outdated. Third, this approach lacks step-by-step reasoning and provides the results in a black-box way.
A natural candidate to overcome the above limitations is to ask the LLM to complete the causal tasks utilizing data input. Two notable examples of such data-enhanced LLMs include i) the Advanced Data Analysis plugin of ChatGPT (OpenAI, 2023a), formerly known as the Code Interpreter (OpenAI, 2023b), which allows users to instruct the LLM through prompts to conduct technical data analysis on self-uploaded datasets; and ii) the latest GPT-3.5/GPT-4 models that include the function calling feature (OpenAI, 2023c), allowing users to introduce new functions for LLM to select from, by providing detailed descriptions and structured information about the function. Additionally, many other recent developments like ToolLLM(Qin et al., 2023b) and GPT4Tools(Yang et al., 2023) also enable LLMs to use external tools to conduct data analysis but are not specifically tailored for the causal tasks.
Despite their ability to analyze external datasets, all of the aforementioned data-enhanced LLMs have the following two limitations in dealing with causal problems: First, in-depth domain knowledge is required to validate the results provided by the current GPT-based models. While these models can provide a wealth of information related to the user’s query, they frequently hallucinate and mislead the user with irrelevant context. For example, they may provide correlation analysis results when causal effect analysis is expected, as shown in Figure 1(a). Second, while some existing approaches may provide guidance for implementing some classical causal methods and suggest the corresponding code script, they fail to provide end-to-end result delivery and easy-to-understand interpretation. Third, pre-trained on massive text corpus, it only provides executions for major methods contained in the training data and requires users’ effort for implementing newly-released approaches as shown in Figure 1(b).
To overcome these challenges, in this study, we develop an end-to-end user-friendly large language model with causal decision-making ability for general audiences. As illustrated in Figure 1, the proposed LLM4Causal model consists of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation. Given a user query and a data file uploaded by the user, the first step transforms the query in natural language to a structured summary in JSON format, which contains essential information such as dataset name, task type, and variable of interest. This transformation is conducted by fine-tuning a pre-trained LLM on the synthetic data, where a well-designed data generation pipeline, is proposed to improve the data quality in terms of both data variety and accuracy. Given the structured JSON data, the second step automatically selects the causal learning algorithm based on the task type and executes it to generate model outputs. The model outputs are then converted to easy-to-understand languages via the third step, where the same LLM model is further fine-tuned on additional data to produce interpretations adhering to three commonly used evaluation rubrics.
Figure 1: A flowchart of the proposed LLM4Causal model consisting of three major steps: user request interpretation, causal tools assignment and execution, and output interpretation.
Our main contributions could be summarized as follows:
  • This paper is the first attempt to construct an end-to-end user-friendly large language model, LLM4Causal, with causal decision-making ability, to the best of our knowledge. Without requiring in-depth domain knowledge, the LLM4Causal could be easily used for general audiences, which addresses the weakness of the current LLM applications on causal tasks.
  • We have finetuned the LLM4Causal to equip it with the capability to conduct end-to-end analysis on causal-related tasks in three steps: i) interpreting user request by converting it to a structured data format, ii) assigning causal tools and executing the corresponding algorithm, and iii) providing easy-to-understand interpretation on the algorithm output.
  • We have proposed a three-step data generation pipeline that combines LLM text generation and human annotation to create customized input-output data. With this well-designed pipeline, we have collected two high-quality benchmark datasets, Causal-Retrieval-Bench for causal function calling and Causal-Interpret-Bench for causal interpretation, with outstanding data variety and accuracy.
  • The proposed LLM4Causal model was extensively evaluated in three major causal decision-making tasks: Causal effect estimation, Causal structure discovery, and Causal policy learning. The proposed method has shown superior performance compared to the benchmark methods while maintaining relatively low training costs.
The rest of the paper is organized as follows. We first introduce the related work in Section 2 and provide the problem formulation in Section 3. In Section 4, we describe the proposed LLM4Causal framework in detail, and the effectiveness of the proposed method is evaluated through extensive experiments in Section 5. Finally, we conclude and discuss some potential future works in Section 6.
(a) User queries on causal relationships.
(b) User queries on implementing LinGAM method.
Figure 2: User interaction with ChatGPT on causal related questions. [^1] The chat histories could be found in link 1, and link 2.
2 Related Work
2.1 Large Language Model with Tool Usage
Despite LLM’s impressive performance in natural language processing tasks, some inherent limitations remain. Examples include the inability to perform precise arithmetic operations (Patel et al., 2021), lack of access to up-to-date information (Komeili et al., 2021), and ignorance of temporal contexts (Dhingra et al., 2022). To overcome these impediments, a natural solution is to enable LLM to utilize tools, such as calculator, search engine, and calendar, to accomplish complex tasks (Qin et al., 2023a). In particular, Schick et al. (2023) proposed Toolformer, a self-supervised language model that could use external tools by calling relevant APIs, without relying on large human-annotated data. With a few tool-usage demonstrations in the instruction prompt, Toolformer is trained to learn which, when, and how APIs could be called to complete a variety of tasks, including question answering, Wikipedia search, and mathematical reasoning. Based on the toolformer architecture, Graph-Toolformer Zhang (2023) is proposed to teach LLMs to use tools on graph reasoning tasks with ChatGPT-augmented prompts (OpenAI, 2022). Moreover, GPT4tools (Yang et al., 2023) are proposed to facilitate LLMs in accomplishing visual-related tasks by generating prompts with multi-modal contexts, and ToolLLM (Qin et al., 2023b) further generalizes the tool-use capability of LLMs to master over 16,000 real-world APIs and enable multiple tool-usage for each query. Focusing on developing frameworks for calling general Application Programming Interfaces (APIs), these methods cannot be directly applied to causal tasks without a well-developed causal API, which currently is not available. On the contrary, our proposed approach is specifically designed for causal tasks based on a framework that is most closely related to Toolformer.
2.2 Enhancing Pre-trained Large Language Models
Pre-trained large language models often contain billions of parameters that impose challenges when fine-tuning the entire model to adapt to downstream tasks. Two classes of research are developed to reduce such a computational burden. A straightforward but effective workaround is freezing most parts of the model and only tuning a small portion of the parameters, which includes additive methods and reparametrization-based methods (Lialin et al., 2023). Additive methods, such as adapters (Houlsby et al., 2019), augment the pre-trained model with extra parameters and only train the newly added parameters. Reparametrization-based methods, such as LoRA (Hu et al., 2021), reduce the number of trainable parameters utilizing low-rank representations. To improve the training efficiency, our proposed method utilizes the LOw-Rank Adaptation method (LoRA) (Hu et al., 2021), which reduces trainable parameters by freezing the pre-trained model and optimizing the rank-decomposition matrices.
Another line of research has demonstrated that template-based prompting can also notably enhance the performance of LLMs. Commonly used prompt types include cloze question style prompts, which consist of string templates with placeholders for the LLM to fill in (Rajagopal et al., 2021; Cui et al., 2021), and prefix style prompts that prepend tokens to the input texts of LLMs (Li and Liang, 2021; Lester et al., 2021). The latter is particularly effective for generative tasks, such as generating interpretations of numeric outputs from causal models in our LLM4Causal framework (Liu et al., 2023). A common way to create such prompts is to design them manually and such prompts usually include a few human-generated exemplars of input-output pairs for LLMs (Chung et al., 2022). These human-crafted prompts have improved performance across various tasks such as question answering and text classification (Brown et al., 2020b; Schick and Schütze, 2020). Recently, Wei et al. (2022) demonstrated that prompting language models to add reasoning steps before the actual output significantly enhances the LLM’s performance on arithmetic, commonsense, and symbolic reasoning tasks. To reduce manual effort in this process, some studies have also explored automation in prompt design. For instance, Shin et al. (2020) discretely optimized the prompt by adding tokens from a certain word collection, while Haviv et al. (2021) focuses on automatic paraphrasing of given manually designed prompts and selecting the most effective one from these variations. In our LLM4Causal framework, we utilize prefix style prompts with manually designed input-output examples to effectively collect data for fine-tuning LLMs to classify user queries into different causal tasks and interpret the numerical results from causal tools. The satisfying performance of our framework has demonstrated the effectiveness of our carefully crafted prompts.
2.3 Causal Decision Making
The decision-making process, crucial in diverse real-life contexts, fundamentally relies on understanding counterfactual scenarios, thereby emphasizing the significance of causal analysis in decision-making. Recently, Song et al. (2023) proposed a comprehensive framework known as Causal Decision Making (CausalDM, Song et al., 2023), which examines various decision-making stages through a causal lens and establishes connections among the relevant literature under the same conceptual umbrella. In general, the body of current literature can be classified into three basic categories: i) causal structure learning (CSL) (Spirtes et al., 2000; Shimizu et al., 2006; Zheng et al., 2018; Yu et al., 2019; Glymour et al., 2019), which focuses on unraveling the complex causal relationships among variables, is pivotal for pinpointing vital factors that influence outcomes, either directly or indirectly; ii) causal effect learning (CEL) (Yao et al., 2021; Hicks and Tingley, 2011), which is primarily concerned with quantifying the effects resulting from different causal paths within a known causal structure, typically concentrating on measuring the difference between outcomes of receiving versus not receiving treatments; and iii) causal policy learning (CPL), which involves evaluating various treatment strategies and developing optimal action plans, frequently investigated as dynamic treatment regime problems (Chakraborty and Murphy, 2014) or conceptualized within the framework of reinforcement learning (Sutton and Barto, 2018; Zeng et al., 2023).
2.4 Large Language Models with Causal Capabilities
The rapid development of LLMs has inspired some innovative methods for causal graph discovery. Kıcıman et al. (2023) is one of the pioneering works in this field, introducing a method that relies solely on "meta-data" (variable names) instead of specific observed variable values. Using LLMs like GPT-3.5 and GPT-4, this approach detects causal relationships between variable pairs through direct querying. Similarly, Long et al. (2023) use GPT-3 to discover causal relationships from variable names within the medical domain, motivated by the capacity of LLM to extract knowledge from the medical texts on which it was trained. However, as Long et al. (2023) highlighted, the efficacy of this type of approach relies heavily on the prompts used for LLM queries and the initial training corpus of the LLM. In a further development, Ban et al. (2023) proposes to integrate the causal relations retrieved directly from LLMs and traditional score-based causal discovery methods. In their framework, the LLM’s causal relations are utilized to guide the optimization process of score-based methods as soft or hard constraints, enhancing the performance compared to using either LLMs or traditional methods alone. Gupta et al. (2023), on the other hand, apply LLMs in selecting causal discovery methods for a specific dataset rather than directly discovering the causal relations on the dataset given. Their framework is claimed to be able to select the causal discovery method with the best F1 score out of a fixed set of candidates. Furthermore, Zhang et al. (2023) discusses the strengths and limitations of LLMs for causal discovery. Although LLMs have advantages in detecting causal relationships within the scope of their training data, their performance suffers when confronted with unfamiliar domains. It is worth mentioning that our approach differs from the previously mentioned studies, as none of them introduces a mechanism to automatically invoke data-based causal discovery methods, nor do they have a mechanism to translate the output into natural languages that general audiences can understand.
3 Problem Formulation
In this paper, we aim to augment existing pre-trained LLMs with the proficiency to address causal inquiries, thereby opening the door to causal decision-making processes for general audiences. As introduced previously, causal decision-making tasks can be classified into three categories: 1) causal structure learning (CSL), 2) causal effect learning (CEL), and 3) causal policy learning (CPL). Each of these primary categories consists of multiple tasks, and our study focuses on five key tasks that are particularly relevant to the interests of our target audience.
Let us denote the user query as $\mathcal{Q}$, the dataset to be analyzed as $\mathcal{D}$, the node set whose interrelationship is of interest as $\mathcal{X}$, the treatment variable as $\mathcal{A}$, the response variable as $\mathcal{Y}$, the mediator variables as $\mathcal{M}$, and the condition of the subgroup of interest as $\mathcal{S}$. The five causal tasks are listed as follows and summarized in Table 1.
Table 1: Constructions of Causal Tasks
Task Dataset Nodes Treatment Response Mediator Condition Function Output Format
($\mathcal{D}$) ($\mathcal{X}$) ($\mathcal{A}$) ($\mathcal{Y}$) ($\mathcal{M}$) ($\mathcal{S}$)
CGL Causal Graph ($\mathcal{G}$)
ATE Numeric Value
HTE Numeric Value
MA Numeric Values
OPO Treatment Level
[CSL] Causal Graph Learning (CGL): Causal graph learning, as its name suggests, involves discovering the causal relationships among variables. This task is crucial for addressing causal queries such as identifying causal links between pairs of variables, determining the variables that influence a particular variable of interest, and quantifying the number of causally connected variable pairs. Given $\mathcal{Q}$ and $\mathcal{D}$, our goal is to learn and report a directed graph, denoted as $\mathcal{G}$, that encapsulates the entire causal structure among variables in $\mathcal{X}$, which is either explicitly specified in $\mathcal{Q}$ or, by default, encompass all variables in $\mathcal{D}$.
[CEL] Average Treatment Effect Estimation (ATE): When the treatment and the target response variable have already been determined, the primary interest lies in quantifying the intervention’s effect size. The objective of the ATE task is to evaluate the average effect of $\mathcal{A}$ on $\mathcal{Y}$ by measuring the difference in counterfactual outcomes between the treated group and the control group. Let $\mathcal{Y}(1)$ denote the response that would be observed if the treatment was applied, and $\mathcal{Y}(0)$ denote the response that would be observed if no treatment is applied. Using the provided dataset $\mathcal{D}$, the treatment variable $\mathcal{A}$, and response variable $\mathcal{Y}$ defined in $\mathcal{Q}$, our objective is to execute appropriate ATE learners to calculate the $\mathbb{E}[\mathcal{Y}(1)-\mathcal{Y}(0)]$.
[CEL] Heterogeneous Treatment Effect Estimation (HTE): Similar to the ATE tasks, the HTE tasks also aim to measure effect sizes. However, HTE tasks diverge by focusing on generating nuanced insights tailored to specific subgroups or individuals, characterized by specific conditions. Using the dataset $\mathcal{D}$, and given the treatment variable $\mathcal{A}$, response variable $\mathcal{Y}$, and specific conditions $\mathcal{S}$ as outlined in $\mathcal{Q}$, we aim to call suitable HTE learners to evaluate the conditionally expected difference $\mathbb{E}[\mathcal{Y}(1)-\mathcal{Y}(0)|\mathcal{S}]$.
[CEL] Mediation Effect Analysis (MA): Mediators, which act as channels for transmitting treatment effects to the response, are prevalent in various real-world applications. For example, an individual’s mood can mediate the relationship between their work environment and physical health, whereas a negative work environment increases stress, leading to a decline in physical health. In order to facilitate better decision-making, it is critical to address the mediators correctly to obtain accurate effect estimates and reveal the complete causal pathway from treatment to outcome. With the provided dataset, and given the $\mathcal{A}$, $\mathcal{Y}$, and $\mathcal{M}$ specified in $\mathcal{Q}$, our objective is to apply appropriate MA estimators to calculate both the direct effect $\mathbb{E}[\mathcal{Y}(1,\mathcal{M}(0))-\mathcal{Y}(0)]$ and the indirect effect $\mathbb{E}[\mathcal{Y}(1)-\mathcal{Y}(1,\mathcal{M}(0))]$. Here, $\mathcal{M}(\cdot)$ is analogously defined to $\mathcal{Y}(\cdot)$, and $\mathcal{Y}(a,\mathcal{M}(a^{\prime}))$ denote the response that would be observed if treatment $a$ was applied and the mediator set to the level that would be observed if treatment $a^{\prime}$ was applied.
[CPL] Off-Policy Optimization (OPO): As the name suggests, OPO task is a one-stop shop for decision-makers. The goal of OPO is to identify the optimal action (i.e., treatment level) that optimizes the expected response to be obtained given the current circumstances. Utilizing the dataset $\mathcal{D}$ provided, with $\mathcal{A}$, $\mathcal{Y}$, and $\mathcal{S}$ explicitly specified in $\mathcal{Q}$, we aim to select suitable policy learners to determine the optimal action $a^{*}={argmax}_{a}\mathbb{E}(\mathcal{Y}(a)|\mathcal{S})$, which would then be recommended to the user. We can provide either single-stage or multi-stage recommendations depending on the data structure of $\mathcal{D}$.
To tackle the aforementioned diverse tasks using a single LLM, we introduce LLM4Causal. By fine-tuning pre-trained LLMs, LLM4Causal is capable of comprehending causal queries, applying appropriate causal tools to analyze the provided tabular dataset, and providing answers by interpreting numerical results in straightforward and fluent language. More technical details of LLM4Causal is discussed in the next section.
4 Proposed Method
In this paper, we introduce a novel three-stage framework to address causality-related tasks, as illustrated in Figure  1. The initial stage involves employing the fine-tuned large language model, LLM4Causal, to interpret users’ inquiries of causal questions by projecting the inquiries into structured templates suitable for functional calls. Specifically, LLM4Causal is carefully fine-tuned with our proposed benchmark dataset, Causal-Retrieval-Bench, enabling it to achieve better performance than other LLMs in classifying causality-related tasks and retrieving relevant information, such as variable names and values. In the second stage, the framework leverages existing tools based on the structured templates generated in the first stage to produce quantitative results for specified causality tasks. However, the numerical outcomes produced in the second stage might not be easily interpretable to a general audience that is unfamiliar with causal inference concepts. To bridge this gap, the third stage of our framework utilizes LLM4Causal to interpret the outputs. To mitigate common pitfalls of LLMs in interpreting numerical outputs, including hallucination and information omissions, the LLM4Causal is further fine-tuned on our proposed Causal-Interpret-Bench dataset. By translating the results from functional calls into clear and easily understandable natural language interpretations, the LLM4Causal addresses the users’ original input queries and provides the interpretation as the final output of the framework.
4.1 STEP1. Interpret User Request
The purpose of the first step is to translate user questions into a causal task classification and a list of related information through two integrated substeps: i) problem classification and ii) attribute extraction. Initially, problem classification categorizes the input question into one of the causal tasks (CGL, ATE, HTE, MA, or OPO) by inferring the underlying user intention. Following this, attribute extraction, conditioned on the task classification, extracts attribute values that are necessary for the expert tools to address the causal problem. Instead of formulating it as a sequential process, we propose to merge them into one sequence-to-sequence (seq2seq) procedure where the output is a structured JSON with the "causal_problem" key and other task-specific keys, e.g. "dataset" ($\mathcal{D}$), "treatment" ($\mathcal{A}$), "outcome" ($\mathcal{Y}$), as keys for the corresponding task. To keep the sequential nature of the problem, we always keep the "causal_problem" key in the first place. The required keys for each causal task are learned implicitly by the parameters of LLM. Some simulated input queries with the corresponding JSON output are shown in Table 2.
Table 2: Example Query Inputs with expected JSON Outputs
Task
Example Input Query
Expected JSON Output
CGL
Does the disaster_risk_reduction.csv dataset provide evidence of a direct link between building code compliance rate and the effectiveness of disaster preparedness campaigns?
{’causal_problem’: [’CSL’, ’CGL’], ’dataset’: [’disaster_risk_reduction.csv’], ’nodes’: [’building_code_compliance_rate’, ’disaster_preparedness_campaigns’]}
ATE
How does the labor force participation rate (labor_participation_rate) in employment.csv contribute to changes in wage growth (wage_increase)?
{’causal_problem’: [’CEL’, ’ATE’], ’dataset’: [’employment.csv’], ’treatment’: [’labor_participation_rate’], ’response’: [’wage_increase’]}
HTE
Based on the findings in the cybersecurity.csv dataset, what impact does the presence of data breach incidents have on cybersecurity investment under a group condition where the cybersecurity readiness index is set at 0.5 (readiness_index=0.5)?
{’causal_problem’: [’CEL’, ’HTE’], ’dataset’: [’cybersecurity.csv’], ’treatment’: [’data_breach_incidents’], ’response’: [’cybersecurity_investment’], ’condition’: [(’readiness_index’, 0.5)]}
MA
Is there substantial evidence in retail.csv indicating that the pathway from retail employment to the e-commerce penetration rate is mediated by the consumer confidence index?
{’causal_problem’: [’CEL’, ’MA’], ’dataset’: [’retail.csv’], ’treatment’: [’retail_employment’], ’response’: [’e-commerce_penetration_rate’], ’mediator’: [’consumer_confidence_index’]}
OPO
If the poverty rate stands at 0.32 (poverty_ratio = 0.32), what recommendations can be derived from the poverty.csv dataset on adjusting social assistance coverage to positively impact the gini coefficient?
{’causal_problem’: [’CPL’, ’OPO’], ’dataset’: [’poverty.csv’], ’treatment’: [’social_assistance_coverage’], ’response’: [’gini_coefficient’], ’condition’: [(’poverty_ratio’, ’0.32’)]}
Using the first pair of examples shown in Table 2 as an illustration, the output JSON has three different pre-set keys conditioned on its related causal task, i.e., Causal Graph Learning (CGL), which is the first object to be inferred.Meanwhile, explicit references to key values are required for the remaining keys, such as "disaster_risk_reduction.csv" for the "dataset" key and [’building_code_compliance_rate’, ’disaster_preparedness_campaigns’] for the "nodes" key.
Even with careful prompt engineering, it is challenging to directly adopt a pre-trained LLM for such a seq2seq task, where the latest GPT4 checkpoint still has a 23% error rate (see Section 5.2.1 for details). Thus, considering the deployment cost and the future potential of adding more features, we fine-tuned a local LLM with an augmented dataset consisting both input queries and output JSONs.
4.1.1 Construct Causal-Retrieval-Bench
To the best of our knowledge, the aforementioned customized input-output pairs can not be collected from the online corpus, since the commonly used LLM training data are mainly conversational materials other than structured data. Therefore, we choose to construct a customized corpus containing both synthetic causal questions and the corresponding JSON representations by prompting an instruction-following LLM.
To collect LLM-generated data, the common approach starts from writing several demonstrated examples. Those demonstrations are then fed to an LLM (e.g., GPT-4) for paraphrasing. We have observed that such an approach suffers from (1) non-controllable topics and variables of interest, which results in content homogeneity, (2) incomplete/non-compliant JSON outputs, and (3) inappropriate paraphrase, where the causal question is transformed to be association-related. To resolve the pain points and prevent data contamination, we propose a three-course procedure, which is illustrated in Figure 3.
Figure 3: Causal-Retrieval-Bench construction procedures for the first step. GPT prompts used in this section can be found in Appendix A. In the left panel, GPT is prompted to list out different topics and measurable variables for each topic. With randomly drawn tasks from section 3 and some variables under the same topic, we generated JSON output in the middle panel. The colored boxes in the left and middle panels are shared topics and variable names of interests. In the right panel, the numbered boxes in template 2 mean blank places to be filled with the task description, the supplied demonstration, JSON inputs, and output restrictions. Common connectors and general instruction in the prompt 2 are shortened into "xxxxx" due to the figure limit.
The procedure can be summarized as an output-first strategy, where the major goal is to improve the correctness and the variety of the output JSON first. An instruction prompt is incorporated to construct causal-related questions in different sentence structures. Specifically, we conducted value generation to prepare a topic database, JSON generation by sampling, and question paraphrasing with prompting.
  • Value Generation: the goal of the first course is to get a candidate pool of the interested topic and variables similar to the RapidAPI [^2] https://rapidapi.com/hub hierarchy tree. We prompt the GPT model to generate some potential topics and related variable names. For each variable, we perform a zero-shot classification by GPT to determine the variable class, i.e., whether it is discrete or continuous. All the collected information is maintained in a three-level hierarchy structure (topic -> variable -> class).
  • JSON Generation: Based on the assumption that causality tasks often involve variables within the same topic, we randomly generate structured JSON data following the topic-variable-class structure. For example, for an average treatment effect estimation task, a topic name is firstly drawn from the stem of the tree as the data file name, and two variables underneath are then considered as candidates for the synthetic treatment and outcome. For causal tasks where a specific group variable is needed, we sampled a random value, either continuous or discrete, based on its class.
  • Question Generation: For each JSON output, we prompt the GPT model to generate five related questions given the context. Utilizing a univariate prompt template, we transform each JSON into an information-augmented prompt consisting of task-specific descriptions, expert-written demo examples, JSON information, and paraphrasing restrictions. Specifically, we implement GPT prompts in this step by a modulized template to improve components’ reusability. Structurally speaking, the prompt starts with a description of the question generation task and the corresponding causal task. It is then followed by information extracted from a JSON from the previous course as the contextual component for the generated question. Subsequently, randomly sampled demonstrations are included as the sample illustrations and output restrictions are itemized as the final remark.
During development, the aforementioned three courses are executed sequentially with quality control on the correctness of generated entities/queries in between. Confused with other statistical concepts, such as correlation and association, the current ChatGPT model has suboptimal performance on the causality-related task. Hence, domain experts are involved after each step to improve the data quality with human annotation. As a result, for the Causal-Retrieval-Bench, we collect 1,500 pairs, i.e., 300 for each causal task, of causal questions with the corresponding JSON outputs.
4.1.2 Fine-tuning a Large Language Model
After collecting a custom dataset containing causal questions and interpreted intents in a structured format, we fine-tune the LLaMA 2 pre-trained model checkpoints for learning underlying language patterns. We denote the input-output pairs as $\mathcal{D}=\{q_{i},j_{i}\}$ for $i=1,2,...,N$, where both inputs $q_{i}$ and outputs $j_{i}=\{c_{i},\mathcal{V}_{i}\}$ can be separated and padded into word tokens $q_{i}=[q_{i1},q_{i2},...,q_{iM}]$ and $j_{i}=[j_{i1},j_{i2},...,j_{iL}]$ with length $M$ and $L$ respectively. A language model checkpoint is denoted as $\mathbb{G}$ which takes the language sequence as input and returns the word probability distribution $P(\cdot)$ for each word for each position with a total length of $L$. The possible input-output word token is limited in a finite dictionary $\mathbb{V}=\{w_{1},w_{2},...,w_{V}\}$ with a cardinal of $V$. We optimize the causal language modeling loss $\ell$, a cross-entropy loss on the token level, with $S$ sample of augmented dataset $\mathcal{S}^{\prime}\subset\mathcal{S}$ as follows:
$\ell(\mathbb{G},\mathcal{S}^{\prime})$ $=$ $\frac{1}{S}\sum_{i=1}^{S}\ell(\mathbb{G}(q_{i}),j_{i})$
$=$ $\frac{1}{SL}\sum_{i=1}^{S}\sum_{k=1}^{L}\textrm{cross-entropy}(\mathbb{G}(q_{i})_{k},j_{ik})$
$=$ $\frac{1}{SL}\sum_{i=1}^{S}\sum_{k=1}^{L}\sum_{l=1}^{V}-I(j_{ik}=w_{l})log(P(\mathbb{G}(q_{i})_{k})=w_{l}),$
where the $\mathbb{G}(q_{i})_{k}=t_{k}$ denotes the output word probability distribution at the $k$-th position, and the cross-entropy loss for that position can be calculated by the logarithm of the probability of word token $j_{ik}$. Minimizing such loss function can encourage the probabilistic model $\mathbb{G}$ to generate similar output to $j_{i}$ when given an input of $q_{i}$.
We implement the Low-Rank Adaptation (LoRA) (Hu et al., 2021) for parameter-efficient fine-tuning since the pre-trained checkpoint has more than $7$ billion parameters and it is computationally inefficient. Instead of changing all parameters in the model, LoRA only tunes adapters’ weight and fixes the entire pre-trained checkpoint. Empirical study shows a relatively small number of ranks like $16$ perform very well in our use case.
4.2 STEP2. Assign and Execute Causal Tools
In this stage, to address user queries, we first select suitable causal learning algorithms based on the task class determined in Step 1, as depicted in Figure 4. Following this selection, the algorithm is executed automatically, using both the extracted information from Step 1 and the user-provided dataset as inputs, to get the estimated result. Depending on the causal task, the output result can be a causal graph $\mathcal{G}$, numerical values of estimated causal effect (estimated expected value of $\mathcal{Y}$), or a treatment level $a\in\mathcal{A}$. LLM4Causal facilitates a variety of approaches for each causal task by leveraging well-known public causal packages, including econml(Oprescu et al., [n. d.]), causalml(Chen et al., 2020), causal-learn(Zheng et al., 2023), and CausalDM(Song et al., 2023). Additionally, we remark that LLM4Causal is designed with flexibility, allowing for the incorporation of new methodologies through the simple integration of function scripts. In the current version, LLM4Causal randomly selects a method for analysis. However, we are in the process of enhancing LLM4Causal for a subsequent release, which aims to provide i) a more interactive platform where users can specify their preferred methods; ii) a comprehensive causal analysis report, offering outputs from all methods with detailed comparisons upon request; and iii) an advanced method selection feature for automatically determining the most suitable approach.
Figure 4: Causal Tool Assignment and Execution in the Second Step
4.3 STEP3. Interpret the Numerical Output
As is discussed in Section 4.2, the function call in our pipeline produces a numerical value of causal effect, a recommended action level, or an adjacency matrix indicating causal relations in a causal graph. Interpreting these outputs needs domain expertise and may not be easily understood by a general audience lacking a causal background. To enhance user-friendliness and reduce the learning curve of LLM4Causal, our final stage entails the interpretation of the function output using fluent, natural language. The process starts with a template interpretation that automatically converts the numerical output to a one-sentence summary (See Appendix B.1 for examples). The final human-readable interpretation is then produced through the utilization of a calibrated LLM, which uses the template summary and other information obtained from previous steps as input and is guided by a designed prompt. The prompt instructs the LLM to avoid generating hallucinatory content, ensure the inclusion of all relevant information, and provide cohesive responses in the relevant context of the original question (See Appendix B.2 for details).
Admittedly, we may directly take the crafted prompt and let a pre-trained LLM such as GPT-4 generate interpretation for us. However, the interpretations obtained this way are problematic. Human annotations indicate that about 25% of the outputs exhibit hallucinations, 10% contain missing information, and 35% lack fluency (see their definition below ). Therefore, we calibrated the same LLM model with additional human annotated interpretation for more accurate result interpretation and the prompt includes the aforementioned entities as basic elements and a graphical illustration can be found in figure 5.
Below, we outline the process used to generate the golden dataset Causal-Interpret-Bench. The approach for model fine-tuning is similar to that used in step 1 and is omitted here for brevity.
Figure 5: Illustration of building Causal-Interpret-Bench in the Third Step. Details about used prompts can be found in Appendix B. In template 3 we fixed the introduction and the final restriction parts because the instruction is about interpretation, and the remaining contextual information is filled differently every time when calling the GPT API. Similar to the idea of section 4.1.1, the general introduction and the restriction on outputs are reused across different causal problems. We feed the original causal query, the causal task associated with the query, and the employed methodology with corresponding function outputs to the prompt as the interpretation context.
4.3.1 Build Causal-Interpret-Bench
Likewise to the procedures in Step 1, we first use the GPT model to generate a silver dataset of interpretations, employing the templates in Appendix B.1 and prompts in Appendix B.2. Among the required inputs, user queries and their corresponding causal task classifications from section 4.1.1 are reused. Using the methodology described in 4.2, we paired the existing information with a causal function. As for the function outputs, the numerical values are randomly generated in formats specific to each task. To create the final golden dataset, we then have annotators manually revise 400 interpretations, adhering to the evaluation rubric outlined below:
  • Hallucination: i) The interpretation incorrectly presents information, such as treating the response variable as the treatment variable and misinterpreting the direct effect in mediation analysis as the total effect; ii) it includes details that are not present in the provided context, such as commenting on the effectiveness of methods used; or iii) it incorrectly applies terms such as ‘correlation’ and ‘association’, which are inappropriate for describing causal relations.
  • Incompleteness: The interpretation lacks one or more essential components, such as the dataset, method, results, and/or all involved variables.
  • Non-fluency: i) The interpretation includes unexplained, meaningless variable names (e.g., ’chg_rate’); ii) it repeats the same content multiple times; or iii) it directly references or rephrases the guidelines in the prompt.
While there are diverse metrics developed to evaluate the factual consistency between the output summaries/interpretation and the original context Zhang et al. (2019) or to assess the conciseness and readability of output summaries, such as the ROUGE score Lin (2004), they fall short of our specific needs. The interpretation we require is highly specialized, necessitating a clear distinction between causality and association, guaranteeing that no crucial information is missed, and ensuring it is comprehensible to general audiences, among other requirements. Current automatic evaluation based on sentence similarity metrics fail to adequately incorporate human preferences, making them unsuitable for our application. Therefore, we carefully defined the three aforementioned rubrics instead to guide the human annotation.
5 Experiments
5.1 End-to-End Performance Evaluation: A Case Study
The user experience of our work is demonstrated in figure 6, 7, and 8, which presents the capabilities of answering causal-related questions with LLM4Causal as a chatbot. For each of the CGL, ATE, and HTE tasks, we show one end-to-end example. The demo website is developed by the Streamlit [^3] https://streamlit.io/ package in Python with the intention that users can chat with our model as a ChatGPT-like environment with causal problem-solving ability. All three examples use a publicly available dataset called "cancer_1000_1.csv" from Li and Beek (2018).
Figure 6: Demo Example of Answering Causal Structure Learning Question
Figure 6 shows the system results (including intermediate steps) for the question "What is the causal relationship between smoking and cancer in the cancer_1000_1.csv dataset?" and LLM4Causal successfully identified it as a CGL task with the JSON output "’causal_problem’: [’CSL’, ’CGL’], ’dataset’: [’cancer_1000_1.csv’], ’covariates’: [None]". The screenshot does not capture them due to the web page limit. After the JSON step (step 1 in 4.1), the service loaded the data from the hard drive and triggered a PC algorithm to discover the causal relationship between the interested variables. Finally, the original questions were answered in natural language, using the results from the graph_learning function.
Figure 7: Demo Example of Answering Average Treatment Effect Question
Figure 8: Demo Example of Answering Heterogeneous Treatment Effect Question
Both figure 7 and figure 8 showcase the full conversation between the user and LLM4Causal. The difference between these two examples is if we wanted the effect estimation limited to a specific group of populations. If so (for example pollution = "high") then we need to trigger the heterogeneous treatment effect instead of the average treatment effect with an additional input "condition". We can see that LLM4Causal discriminated different problems prosperously and provided reasonable interpretations based on the provided context.
5.2 Ablation Analysis
We also carefully evaluated the model performance with an ablation analysis to determine the effectiveness of the causal entity extraction described in section 4.1. Our fine-tuned model is benchmarked with a GPT4 variant (with or without function call features). As a further step, more simulation studies involving end-to-end and interpretation evaluation will be released soon.
As for LLM4Causal, we trained two different checkpoints: ‘LLM4Causal-Retrieve’ only trains on Causal-Retrieval-Benchand ‘LLM4Causal-Mixed’ or ‘LLM4Causal’ simultaneously trained on both datasets. The purpose of showcasing ‘LLM4Causal-Mixed’ and ‘LLM4Causal-Retrieve’ is to show that mixing both sources of the augmented data will not harm the inference quality and that only having one univariate model is sufficient.
5.2.1 Causal Entity Extraction
We left 150 samples of the training data from the process of section 4.1. For each simulated causal question, the benchmarked model was asked to produce a JSON output following the structure rule for the downstream causal function, and the output was validated by the ground truth label by comparing the value for each key. The accuracy scores for correctly specifying the causal task classification and exactly matching the values under each key are reported. As a consequence of different causal tasks having variant keys, if the model misclassifies the causal task in the first place, its performance for the corresponding keys to the problem will be harmed.
Here we are benchmarking GPT4 with the function call feature enabled. It was developed by OpenAI in 2023 and demonstrated a strong capability of zero-shot function selection and argument extraction. Multiple papers (Qin et al., 2023b; Zhang, 2023) treated its output as the golden label for local LLM development. We created one description file for each causal task with the specification of required input arguments, the output was mapped into one JSON by with an additional "causal_problem" key storing the selected function.
Table 3: Causal Entity Extraction Performance
Causal Task Dataset Nodes Treatment Response Mediator Condition
gpt3.5-turbo 0.77 0.77 0.27 0.60 0.61 0.81 0.33
gpt3.5-turbo (guided) - 1.00 0.27 0.84 0.80 0.94 0.42
gpt4-turbo 0.77 0.77 0.27 0.67 0.62 0.67 0.32
gpt4-turbo (guided) - 1.00 0.27 0.90 0.86 0.97 0.40
LLM4Causal-Retrieve 0.98 0.95 0.81 0.97 0.98 0.94 0.97
LLM4Causal-Mixed 0.99 0.99 0.88 0.98 0.97 1.00 0.98
It is shown in the table 3 that the turbo version of both GPT3.5 and GPT4 can perform the causal task identification task relatively well, but their accuracy on detailed argument extraction is limited, especially for keys with complicated formats such as condition and nodes. As a further step, we only provided the correct function to the model (marked as guided) and tried to investigate if the performance was masked by selecting the incorrect causal task. We can see significant improvement as expected. For gpt4 specifically, we had an increase in the dataset ($0.77\rightarrow 1.00$), treatment ($0.67\rightarrow 0.90$), response ($0.62\rightarrow 0.86$), and mediator ($0.67\rightarrow 0.97$), but it still made considerable mistakes on exactly matching nodes and conditions. As for LLM4Causal, we can see that both checkpoints can solve such a seq2seq problem very well, only the Condition key having an exact match accuracy rate that is less than $0.9$. A seemingly surprising fact is the mixed version even outperforms the retrieve-only version on most tasks except the nodes, which may be due to the inclusion of the interpretation data, which also involves causal queries and their corresponding causal tasks with interpretations, increasing the word connection between causal tasks and their meanings.
6 Conclusion
In this paper, we have proposed LLM4Causal, the first end-to-end user-friendly large language model with causal decision-making ability. Without requiring in-depth domain knowledge, the proposed model could be easily used for the general audience, which remedies the weakness of the current LLM application on the causal tasks. Furthermore, we have proposed a novel data generation pipeline by utilizing GPT-based LLM models and human annotations to improve the data quality in terms of both data variety and accuracy. Fine-tuned to interpret users’ query and function outputs, LLM4Causal has shown superior performance in causal entity extraction. As a further step, more simulation studies involving end-to-end and interpretation evaluation will be released soon.
To conclude, we briefly discuss some limitations followed by inspired possible future directions. First, the proposed model is only capable of handling five major causal tasks, but it is worth mentioning that our framework could be easily extended to support more causal tasks with cutting-edge methodologies. Second, the current approach randomly selects one causal tool in step 2, but it could be a valuable direction to enable the model to automatically select the best available tools based on data characteristics and context information. Third, in this project, we exploit the LLM’s capability of using statistical tools to conduct data analysis, while the internal knowledge that LLM obtains during its training may also be beneficial. It could be of great value to integrate LLM’s tool-usage ability with LLM’s internal knowledge to further improve its performance on causal tasks. Lastly, a potential direction to explore is the interactive potential of the framework so that the proposed method could iteratively adapt to users’ feedback, which could be beneficial in following the users’ preferences for causal tools, encouraging users to collect necessary information when their input is not enough, and enhancing the interpretations based on their extra input in the problem context.
Appendix A Details about Step 1
The overall cloze template contains 5 different missing pieces. Among them, the task requirement is a verbal description of the task like the role definition of ChatGPT. JSON information is followed by listing out each item line by line. Further, some randomly drawn human-written samples, split into two parts, are added as demonstrations for the GPT model to further understand what kind of output we need. Finally, we add the requirements for the output in the last part for better control of the GPT output.
Given the details below, generate five diverse and plain-worded questions [Task requirements]:
- [JSON information]
Here are some examples:
Example 1: given the following information:
- [Demonstration 1 part A]
The corresponding questions can be:
- [Demonstration 1 part B]
[Requirements]
Below is an example of the full instruction for Causal Structure Learning data generation with information filled in:
Given the details below, generate five diverse and plain-worded questions asking about the existence
of causal effects between interested variables:
- Dataset name: social_media.csv
- Interested variable: number of social media users, user demographics
Here are some examples:
Example 1: given the following information:
- Dataset name: employee_data.csv
- Interested variable: all variables
The corresponding questions can be:
- Is there a method to discover every direct influence present in the employee_data.csv dataset?
- Within the employee_data.csv dataset, how many instances of one factor directly causing another
can be observed?
- How many causal connections can be identified in the employee_data.csv dataset?
Example 2: given the following information:
- Dataset name: job_data.csv
- Interested variable: education (edu), job satisfaction (job_satisfaction)
The corresponding questions can be:
- Within the job_data.csv dataset, is there causal links between education (edu) and job satisfaction
(job_satisfaction)?
- Does the job_data.csv data reveal any direct relationships between education (edu) and job
satisfaction (job_satisfaction)?
- Are there discernible causal links between education (edu) and job satisfaction (job_satisfaction)
in the job_data.csv dataset?
Example 3: given the following information:
- Dataset name: weather_data.csv
- Interested variable: temperature (temp), humidity (humid), rainfall (rain), local latitude (lat),
local altitude (alt)
The corresponding questions can be:
- What are causal links among temperature (temp), humidity (humid), rainfall (rain), local latitude
(lat), and local altitude (alt) in dataset weather_data.csv?
- Within the dataset weather_data.csv, what is the nature of the causal links among temperature (temp),
humidity (humid), rainfall (rain), local latitude (lat), and local altitude (alt)?
- In the weather_data.csv dataset, what causal connections can be identified among temperature (temp),
humidity (humid), rainfall (rain), local latitude (lat), and local altitude (alt)?
Ensure that the questions employ a mix of different phrasing and diverse sentence structures.
Ensure that the questions do not mention correlation and association.
Ensure that the questions place a strong emphasis on the effect existence or the total number.
Ensure that all variable names under interested variables field exist in all five quetsion.
Ensure that the provided interested variable names are integrated naturally into the questions without
discarding or altering any part of them.
Appendix B Details about Step 3
B.1 Template Outcome
The one-sentence summary of the function outcome uses natural language to express the same meaning. The benefit of using such templates is to aid LLM in understanding this information, so it can better utilize the sentence (compared with just a number) to answer the initial question.
For the CSL task where the total outcome is an adjacency matrix of variables of interest, we picked the top K causal relationships defined by the significance of the independence test. The summary would contain two sentences, the first sentence is "There are {K} pairs of significant causal relationships." where K is the number of detected (equal or less than 2) causal links. In the second sentence, we list out these top causal paths by the prompt "The {X} would causally influence the {Y}" where X and Y are two variable names of interest.
As for the ATE and the HTE tasks, we expressed the learned effect size by "The average treatment effect of setting {A} as 1 on the {Y} is {effect size}." and "The heterogeneous treatment effect of setting {A} as 1 on the {Y} is {effect size} for those having {X} = {value}". Here A, Y, and X denote the treatment, outcome, and condition variable names separately. The effect size is the learned effect size for either ATE or HTE, and the value in HTE is the covariate condition.
The Mediation analysis (MA) task has three numerical outputs: a total effect indicating the total scale of the causal effect from a treatment variable A to the outcome Y, a mediation effect that goes through mediation variable M (A->M->Y), and a direct effect only from A->Y. Two separate sentence prompts, "The overall impact of the {A} on the {Y} is {total effect}." and "This comprises a direct effect of {direct effect} from the {A} to the {Y}, and an indirect effect of {mediation effect}, mediated by the {M}."
Finally, for CPL we used the prompt "The best action of the {A} is {A} = {a recommended value}." because here we focus on the policy learning task and a recommended treatment given a condition is the function outcome.
One example prompt filled with variable names of each causal task is provided in the code block below.
CSL: "There are 3 pairs of significant causal relationships. The gender_index would causally
influence the diversity_index. The gender_index would causally influence the LGBTQ_inclusion.
The disability_inclusion_index would causally influence the LGBTQ_inclusion."
ATE: "The average treatment effect of setting homeownership_rate as 1 on the affordability_index
is 0.45."
HTE: "The heterogeneous treatment effect of setting professional_athlete_salaries as 1 on the
event_attendance is -1.41 for those having medal_tally = 0.79."
MA: "The overall impact of the age_distribution on the gender_ratio is 16.17. This comprises a
direct effect of 9.43 from the age_distribution to the gender_ratioand an indirect effect of 6.74,
mediated by the migration_speed."
CPL: "The best action of the professional_athlete_salaries is professional_athlete_salaries = C."
B.2 Prompt
The instruction prompt for GPT and other LLMs to generate interpretation is as follows, which lists out requirements stated in the evaluation criteria 4.3.1:
(A) is a list of information that includes i) the original causal problem, ii) the class
identification of the causal problem, iii) the used method, and iv) the outcomes.
Interpret the results in (A) in response to the original causal problem, using neutral
language to paraphrase it more fluently and engagingly.
The output summary is (I)
Guidelines:
1: (I) must concentrate on interpreting the result provided in (A) in response to the problem.
2: (I) must include all the results, methods, and dataset name in (A).
3: (I) may include jargon from (A), but it shouldnot include any other technical terms
not mentioned in (A).
4: The problem in (A) is a causal problem, thus (I) should not interpret the results as
correlation or association.
5: (I) should use a diversified sentence structure that is also reader-friendly and
concise, rather than listing information one by one.
6: Instead of including the problems, (I) should use the original problem to develop a
more informative interpretation of the result.
7: (I) has to avoid using strong qualifiers such as ’significant’.
8: (I) has to be {n_sentences} sentences or less long, with no repetition of contents.
9: (I) must not comment on the results.
(A):
i) original causal problem: {query}
ii) class identification of the causal problem: {problem}
iii) used method: {method}
iv) outcomes: {function_out}
(I):